x +0.4x + 2x = 1 - Hz(t)x(0) =0, x0) = 0solve using laplace transforms
Answer #1

CA V +0.4x +2x = 1 - Hş (+) Taking Laplace Transformation Worit. t. [] +0.4 d [x] +2 d [x] = { [t] - d [H5 (4)] det, [x(b7 x (t) = f(t) – H5 Ct) f The f6)-1(10:] 2 Rc1 2 25 5 +2 3) 1 = A (3+ x + 2) + (B5+C)s. Comparing the co-efficients of 1,5, 52 Ś 10 • 3-4 (4746[***--*-043-45). ferte e o 494) - to ņē simu 569 = 4- e *((*) - tip mo siya ng The Solution of the ODE , *Feel free to ask any doubt in comment section. Thank you sir/ma'am. ??